# Graphical Calculus for Non-Gaussian Quantum States

### v2

@@ Abstract
We provide a graphical method to describe and analyze non-Gaussian quantum states using a hypergraph framework. These states are pivotal resources for quantum computing, communication, and metrology, but their characterization is hindered by their complex high-order correlations. The formalism encapsulates transformation rules for any Gaussian unitary operation and local quadrature measurement, offering a visually intuitive tool for manipulating such states through experimentally feasible pathways. Notably, we develop methods for the generation of complex hypergraph states with more or higher-order hyperedges from simple structures through Gaussian operations only, facilitated by our graphical rules. We present illustrative examples on the preparation of non-Gaussian states rooted in these graph-based formalisms, revealing their potential to advance continuous-variable general quantum computing capabilities.
@@ 1 Introduction
In the quest to demonstrate a genuine quantum advantage in the continuous variable (CV) domain, substantial progress on the controllable generation of multimode non-Gaussian states has recently been made by photon subtraction [1, 2, 3] and nonlinear unitary operations such as spontaneous parametric down-conversion (SPDC) [4]. These strategies usually introduce higher-order statistical moments of the quadrature field operators, go beyond Gaussian states and therefore have irreplaceable advantages as quantum resource [5, 6] in various tasks, including entanglement distillation [7, 8], quantum metrology [9], and CV quantum computation [10, 11, 12, 13]. Compared to the remarkable progress in experimental realizations, how to understand and more definitely classify the multimode non-Gaussian states with high-order interplay among quadratures, still faces a formidable challenge. To address it in a more intuitive and systematic manner, graphical formalisms [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] as a means of describing quantum states emerges as a highly promising solution.
Figure 1: Examples of graph and hypergraph states, as well as the graphical calculus for hypergraph states as discussed in this paper. (a) Topology of a four-mode square graph state and the circuit to generate it. (b) A four-mode hypergraph state. Hypergraph states generalize graph states by allowing hyperedges (colored closed shapes), which represent interactions between multiple modes. (c) Based on the graphical rules proposed in this paper, a complex hypergraph state with more or higher-order hyperedges can be generated from simple structures solely through Gaussian operations.
@@
Here, we provide a unified graphical calculus that serves as a comprehensive framework for manipulating non-Gaussian states in CV systems. This calculus encapsulates hypergraph transformation rules encompassing all single-mode and typical two-mode Gaussian unitary operations, together with local quadrature measurements, offering an intuitive visual modality for implementing experimentally feasible operations on CV non-Gaussian states. Crucially, as shown in Fig. 1(c), we demonstrate the capability to generate complex hypergraph states with more or higher-order hyperedges from simple structures solely through Gaussian operations, underscoring the power and versatility of the graphical rules provided here. Furthermore, we give three examples of non-Gaussian state preparation schemes rooted in our graph-based formalism. These instances illustrate the practical potential of the graphical calculus, positioning it as a valuable tool for advancing the frontiers of quantum information processing.
@@ 2 CV Hypergraph States
A weighted hypergraph is a tuple $H=(V,E,T)$ with a set $V=\{v_{i}\}_{i=1}^{n}$ of $n$ vertices, a set $E=\{e_{j}\}_{j=1}^{|E|}$ of ${|E|}$ hyperedges, and a set $T=\{t_{e_{j}}\}_{j=1}^{|E|}$ of real valued weights associated to hyperedges, where $|E|$ denotes the number of elements of $E$. A CV hypergraph state |$\ket{H}$ is defined as an $n$-mode quantum state where the vertices represent the modes of the state and the weighted hyperedges represent generalized controlled-$Z$ gates performed on the modes [40, 39]. More formally, we have |$\ket{H}=\prod_{e_{j}\in E}C_{e_{j}}(t_{e_{j}})\ket{0}_{p}^{\otimes n}$, where |$\ket{0}_{p}$ is the zero-momentum eigenstate in the limit of infinite squeezing with normalization $\bra{p^{\prime}}\ket{p}=\delta_{p,p^{\prime}}$ and $C_{e_{j}}(t_{e_{j}})=\operatorname{e}^{it_{e_{j}}\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\dots\operatorname{\mathit{\hat{q}}}_{v_{m}}}$ is a generalized controlled-$Z$ gate acting on $m$ modes connected by the $j-$th hyperedge $e_{j}=\{v_{1},v_{2},\dots v_{m}\}$. If a hyperedge only contains two vertices, $C_{e}$ is the standard controlled-$Z$ gate $C_{e}(t)=\operatorname{e}^{it\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}}$ and if it contains three vertices, it is the Toffoli gate $C_{e}(t)=\operatorname{e}^{it\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\operatorname{\mathit{\hat{q}}}_{v_{3}}}$ [41]. We use the shorthand notation $q_{e_{j}}\coloneqq\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\dots\operatorname{\mathit{\hat{q}}}_{v_{m}}$. A hypergraph state of $k$-th order is classified by the maximum number of $k$ modes connected with one hyperedge.
@@
In order to present the graphical calculus more conveniently, we first introduce some basic rules of set operations. The hyperedge set $E_{a}$ is defined as the set of all hyperedges containing vertex $a$. The adjacency set $\operatorname{\mathcal{A}}(a)$ of vertex $a\in V$ is defined as $\operatorname{\mathcal{A}}(a)=\{e\setminus\{a\}\mid e\in E_{a}\}$. Considering the bijective mapping between hyperedge and weight sets, we define the union $(E_{U};T_{U})=(E_{1};T_{1})\cup(E_{2};T_{2})$ as a union of all tuples. For hyperedges appearing in both sets, we add up their weights. For the hypergraph in Fig. 2(a), the set of hyperedges containing mode $A$ is given by $E_{A}=\{\{A,B,C\},\{A,D\}\}$ with associated weights $T_{A}=\{1,2\}$ and $\mathcal{A}(A)=\{\{B,C\},\{D\}\}$. If we unite $E_{A}$ and $E_{B}=\{\{A,B,C\}\}$ with $T_{B}=\{1\}$, $(E_{U};T_{U})=(\{\{A,B,C\},\{A,D\}\};\{2,2\})$ is obtained.
Although not all quantum states can be transformed under the graphical rules, some states are able to be represented indirectly in the form of hypergraphs, for example, the state after applying the cubic phase gate $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}^{3}}$ which contains a third-order moment term of one single-mode operator and plays important roles in universal quantum computing [42, 11, 43]. We will demonstrate later that it can be represented as a sequence of operations similar to those of our formalism.
@@ 3 Results
In this section, we provide unified graphical rules of Gaussian unitary operations and Gaussian measurements for CV hypergraph states. We show the rules of single-mode Gaussian operations including displacement, shearing, squeezing, and rotation [44, 45] and further discuss the CV controlled-$Z$ gate. Derivations are given in Appendix A. As any multimode Gaussian operation can be represented as a concatenation of a beam splitter, single-mode squeezing, and phase shifts through the Bloch-Messiah decomposition [46], where the beam-splitter operation is equivalent to a controlled-$Z$ gate [41], any multimode Gaussian operation can be represented in our graphical calculus.
@@ 3.1 Displacement
However, when performing the position displacement $X^{a}(s)=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}$ to the hypergraph state, due to the incompatibility between two operators $\operatorname{\mathit{\hat{q}}}_{a}$ and $\operatorname{\mathit{\hat{p}}}_{a}$, additional correlation terms will be generated when they simultaneously act on the zero-momentum eigenstate |$\ket{0}_{p}$. It influences hyperedges of the adjacency $\mathcal{A}(a)$ of mode $a$ such that $(E^{\prime},T^{\prime})=(E,T)\cup(\operatorname{\mathcal{A}}(a),\{\tilde{t}\})$, where $\tilde{t}\coloneqq-s\times t_{e\cup a}$, which means the position displacement on mode $a$ will add or change hyperedges between its adjacent modes with the weight multiplied by $-s$. For example, when the position displacement $X^{A}(1)$ is applied to mode $A$ of the hypergraph state corresponding to Fig. 2(a), two new hyperedges between vertices $\{B,C\}$ and $\{D\}$ of $\operatorname{\mathcal{A}}(a)$ are created, as represented by the purple line and circle in Fig. 2(c).
@@ 3.2 Shearing
The shearing operator includes the momentum shearing $D_{p}^{a}(s)=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{p}}}_{a}^{2}}$ and the position shearing $D_{q}^{a}(s)=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{q}}}_{a}^{2}}$, which shears the state with respect to the $\operatorname{\mathit{\hat{p}}}$ or $\operatorname{\mathit{\hat{q}}}$-axis by a gradient of $s$. The shearing operator introduces an additional second-order moment term in the exponent, which does not occur in a standard hypergraph state, so this operation leaves the set of hypergraph states. But this is in line with observations made on hypergraph states with qubits [48]. For the position shearing operation, a term without effecting the hyperedge set of the state, $D_{q}^{a}(s)\ket{H}=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{q}}}_{a}^{2}}\ket{H}$, will be created. For example, performing the position shearing $D_{q}^{B}(2)$ on mode $B$ will change the hypergraph into the one shown in Fig. 2(d). The dotted line indicates that there is a second-order moment on its own.
Similarly, due to the incompatibility between two operators $\operatorname{\mathit{\hat{q}}}_{a}$ and $\operatorname{\mathit{\hat{p}}}_{a}$, the momentum shearing operation $D_{p}^{a}(s)$ on mode $a$ will introduce new terms acting on the adjacency $\operatorname{\mathcal{A}}(a)$. More precisely, we have $D_{p}^{a}(s)\ket{H}=e^{i\frac{s}{2}(\sum_{e\in\operatorname{\mathcal{A}}(a)}t_{e\cup a}\operatorname{\mathit{\hat{q}}}_{e})^{2}}\ket{H}$. In most cases, momentum shearing operation leads to a state which is not a hypergraph state. In the example of Fig. 2(a), the operation $D_{p}^{A}(1)$ on mode $A$ will introduce the term $\operatorname{e}^{i(\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}+2\operatorname{\mathit{\hat{q}}}_{D})^{2}/2}$ to the original hypergraph state. Expanding the square term we get a forth-order moment term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}/2}$, a second-order moment term $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{D}^{2}}$, and a hyperedge $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}}$ among modes $\{B,C,D\}$, as shown in Fig. 2(e).
Therefore, in a hypergraph state of $k$-th order, a new hyperedge of $k$-th order can be generated to connect different modes by shearing operations. This effect can also link more pairs in the graph state. In principle, if a hypergraph state contains one $k$-th order hyperedge and another $l$-th order hyperedge which intersect in exactly one mode, shearing operations acting on this shared mode can expand the maximum order to $(k+l-2)$-th, offering a practical approach to generate more or higher-order hypergraph states with multimode nonlinearity using only local Gaussian operations.
@@ 3.3 Squeezing
Applying a single-mode squeezing operator $S^{a}(s)=\operatorname{e}^{-i\frac{s}{2}(\operatorname{\mathit{\hat{p}}}_{a}\operatorname{\mathit{\hat{q}}}_{a}+\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a})}$ on mode $a$ multiplies the weight of all hyperedges linked to mode $a$ by an additional factor of $\operatorname{e}^{-s}$, while other hyperedges remain unchanged. In Fig. 2(f), we show an example by squeezing mode $A$ with $S^{A}(1)$, hence, $t_{AD}=2\operatorname{e}^{-1},t_{ABC}=\operatorname{e}^{-1}$. Therefore, by using a squeezing operation, the value of weights in the hyperedges can be effectively proportionally adjusted, that is, the strength of the correlation between different modes can be changed.

### v3

# Graphical Framework for Non-Gaussian Quantum States

@@ Abstract
We provide a graphical method to describe and analyze non-Gaussian quantum states using a hypergraph framework. These states are pivotal resources for quantum computing, communication, and metrology, but their characterization is hindered by their complex high-order correlations. The framework encapsulates transformation rules for a series of typical Gaussian unitary operation and local quadrature measurement, offering a visually intuitive tool for manipulating such states through experimentally feasible pathways. Notably, we develop methods for the generation of complex hypergraph states with more or higher-order hyperedges from simple structures through Gaussian operations only, facilitated by our graphical rules. We present illustrative examples on the preparation of non-Gaussian states rooted in these graph-based formalisms, revealing their potential to advance continuous-variable general quantum computing capabilities.
@@ 1 Introduction
Figure 1: Examples of graph and hypergraph states, as well as the graphical framework for hypergraph states as discussed in this paper. (a) Topology of a four-mode square graph state and the circuit to generate it. (b) A four-mode hypergraph state. Hypergraph states generalize graph states by allowing hyperedges (colored closed shapes), which represent interactions between multiple modes. (c) Based on the graphical rules proposed in this paper, a complex hypergraph state with more or higher-order hyperedges can be generated from simple structures solely through Gaussian operations.
In the quest to demonstrate a genuine quantum advantage in the continuous variable (CV) domain, substantial progress on the controllable generation of multimode non-Gaussian states has recently been made by photon subtraction [1, 2, 3] and nonlinear unitary operations such as spontaneous parametric down-conversion (SPDC) [4]. These strategies usually introduce higher-order statistical moments of the quadrature field operators, go beyond Gaussian states and therefore have irreplaceable advantages as quantum resource [5, 6] in various tasks, including entanglement distillation [7, 8], quantum metrology [9], and CV quantum computation [10, 11, 12, 13]. Compared to the remarkable progress in experimental realizations, how to understand and more definitely classify the multimode non-Gaussian states with high-order interplay among quadratures, still faces a formidable challenge. To address it in a more intuitive and systematic manner, graphical frameworks [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] as means of describing quantum states emerge as a highly promising solutions.
@@
Here, we provide graphical rules which serve as a comprehensive framework for manipulating non-Gaussian states in CV systems. This pictorial representation encapsulates hypergraph transformation rules encompassing typical Gaussian unitary operations, together with local quadrature measurements, offering an intuitive visual modality for implementing experimentally feasible operations on CV non-Gaussian states. Crucially, as shown in Fig. 1(c), we demonstrate the capability to generate complex hypergraph states with more or higher-order hyperedges from simple structures solely through Gaussian operations, underscoring the power and versatility of the graphical rules provided here. Furthermore, we give three examples of non-Gaussian state preparation schemes rooted in our graph-based formalism. These instances illustrate the practical potential of the graphical framework, positioning it as a valuable tool for advancing the frontiers of quantum information processing.
@@ 2 CV Hypergraph States
A weighted hypergraph is a tuple $H=(V,E,T)$ with a set $V=\{v_{i}\}_{i=1}^{n}$ of $n$ vertices, a set $E=\{e_{j}\}_{j=1}^{|E|}$ of ${|E|}$ hyperedges, and a set $T=\{t_{e_{j}}\}_{j=1}^{|E|}$ of real valued weights associated to hyperedges, where $|E|$ denotes the number of elements of $E$. A CV hypergraph state |$\ket{H}$ is defined as an $n$-mode quantum state where the vertices represent the modes of the state and the weighted hyperedges represent generalized controlled-$Z$ gates performed on the modes [44, 43]. More formally, we have
|$\ket{H}=\prod_{e_{j}\in E}C_{e_{j}}(t_{e_{j}})\ket{0}_{p}^{\otimes n},$ (1)
where |$\ket{0}_{p}$ is the zero-momentum eigenstate in the limit of infinite squeezing with normalization $\bra{p^{\prime}}\ket{p}=\delta_{p,p^{\prime}}$ and $C_{e_{j}}(t_{e_{j}})=\operatorname{e}^{it_{e_{j}}\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\dots\operatorname{\mathit{\hat{q}}}_{v_{m}}}$ is a generalized controlled-$Z$ gate acting on $m$ modes connected by the $j-$th hyperedge $e_{j}=\{v_{1},v_{2},\dots v_{m}\}$. If a hyperedge only contains two vertices, $C_{e}$ is the standard controlled-$Z$ gate $C_{e}(t)=\operatorname{e}^{it\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}}$ and if it contains three vertices, it is the Toffoli gate $C_{e}(t)=\operatorname{e}^{it\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\operatorname{\mathit{\hat{q}}}_{v_{3}}}$ [45]. We use the shorthand notation $q_{e_{j}}\coloneqq\operatorname{\mathit{\hat{q}}}_{v_{1}}\operatorname{\mathit{\hat{q}}}_{v_{2}}\dots\operatorname{\mathit{\hat{q}}}_{v_{m}}$ and define the quadrature operators in units such that $[\operatorname{\mathit{\hat{q}}},\operatorname{\mathit{\hat{p}}}]=i$ holds. A hypergraph state of $k$-th order is classified by the maximum number of $k$ modes connected with one hyperedge.
@@
In order to present the graphical framework more conveniently, we first introduce some basic rules of set operations. The hyperedge set $E_{a}$ is defined as the set of all hyperedges containing vertex $a$. The adjacency set $\operatorname{\mathcal{A}}(a)$ of vertex $a\in V$ is defined as $\operatorname{\mathcal{A}}(a)=\{e\setminus\{a\}\mid e\in E_{a}\}$. Considering the bijective mapping between hyperedge and weight sets, we define the union $(E_{U};T_{U})=(E_{1};T_{1})\cup(E_{2};T_{2})$ as a union of all tuples. For hyperedges appearing in both sets, we add up their weights. For the hypergraph in Fig. 2(a), the set of hyperedges containing mode $A$ is given by $E_{A}=\{\{A,B,C\},\{A,D\}\}$ with associated weights $T_{A}=\{1,2\}$ and $\mathcal{A}(A)=\{\{B,C\},\{D\}\}$. If we unite $E_{A}$ and $E_{B}=\{\{A,B,C\}\}$ with $T_{B}=\{1\}$, $(E_{U};T_{U})=(\{\{A,B,C\},\{A,D\}\};\{2,2\})$ is obtained.
Although not all quantum states can be transformed under the graphical rules, some states are able to be represented indirectly in the form of hypergraphs. For example, the state after applying the cubic phase gate $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}^{3}}$ which contains a third-order moment term of one single-mode operator and which plays important roles in universal quantum computing [46, 11, 47]. We will demonstrate later that it can be represented as a sequence of operations similar to those of our formalism.
@@ 3 Results
In this section, we provide unified graphical rules of Gaussian unitary operations and Gaussian measurements for CV hypergraph states. We show the rules of single-mode Gaussian operations including displacement, shearing, squeezing, and rotation [48, 49] and further discuss the CV controlled-$Z$ gate. Derivations are given in App. A. In principle, any multimode Gaussian operation can be represented as a concatenation of a beam splitter, single-mode squeezing, and phase shifts through the Bloch-Messiah decomposition [50], where the beam-splitter operation is equivalent to a controlled-$Z$ gate [45]. Note that some of the fundamental operations generate additional terms, which need to be taken into account when applying multiple operators consecutively.
@@ 3.1 Displacement
However, when performing the position displacement $X^{a}(s)=\operatorname{e}^{-is\operatorname{\mathit{\hat{p}}}_{a}}$ to the hypergraph state, due to the incompatibility between two operators $\operatorname{\mathit{\hat{q}}}_{a}$ and $\operatorname{\mathit{\hat{p}}}_{a}$, additional correlation terms will be generated when they simultaneously act on the zero-momentum eigenstate |$\ket{0}_{p}$. It influences hyperedges of the adjacency $\mathcal{A}(a)$ of mode $a$ such that $(E^{\prime};T^{\prime})=(E;T)\cup(\operatorname{\mathcal{A}}(a);\{\tilde{t}\})$, where $\tilde{t}\coloneqq-s\times t_{e\cup a}$. This means that the position displacement on mode $a$ will add or change hyperedges between its adjacent modes with the weight multiplied by $-s$. For example, when the position displacement $X^{A}(1)$ is applied to mode $A$ of the hypergraph state corresponding to Fig. 2(a), two new hyperedges between vertices $\{B,C\}$ and $\{D\}$ of $\operatorname{\mathcal{A}}(a)$ are created, as represented by the purple line and circle in Fig. 2(c).
@@ 3.2 Shearing
The shearing operator includes the momentum shearing $D_{p}^{a}(s)=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{p}}}_{a}^{2}}$ and the position shearing $D_{q}^{a}(s)=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{q}}}_{a}^{2}}$, which shears the state with respect to the $\operatorname{\mathit{\hat{p}}}$ or $\operatorname{\mathit{\hat{q}}}$-axis by a gradient of $s$. The shearing operator introduces an additional second-order moment term in the exponent, which does not occur in a standard hypergraph state, so this operation leaves the set of hypergraph states. This is in line with observations made on hypergraph states with qubits [52]. For the position shearing operation, a term without effecting the hyperedge set of the state, $D_{q}^{a}(s)\ket{H}=\operatorname{e}^{i\frac{s}{2}\operatorname{\mathit{\hat{q}}}_{a}^{2}}\ket{H}$, will be created. For example, performing the position shearing $D_{q}^{B}(2)$ on mode $B$ will change the hypergraph into the one shown in Fig. 2(d). The dotted line indicates that there is a second-order moment on its own. We discuss how to deal with these extra decorations in App. A.
Similarly, due to the incompatibility between two operators $\operatorname{\mathit{\hat{q}}}_{a}$ and $\operatorname{\mathit{\hat{p}}}_{a}$, the momentum shearing operation $D_{p}^{a}(s)$ on mode $a$ will introduce new terms acting on the adjacency $\operatorname{\mathcal{A}}(a)$. More precisely, we have $D_{p}^{a}(s)\ket{H}=e^{i\frac{s}{2}(\sum_{e\in\operatorname{\mathcal{A}}(a)}t_{e\cup a}\operatorname{\mathit{\hat{q}}}_{e})^{2}}\ket{H}$. In most cases, momentum shearing operation leads to a state which is not a hypergraph state. In the example of Fig. 2(a), the operation $D_{p}^{A}(1)$ on mode $A$ will introduce the remaining unitary $\operatorname{e}^{i(\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}+2\operatorname{\mathit{\hat{q}}}_{D})^{2}/2}$ to the original hypergraph state. Expanding the square term we get a forth-order moment term $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}/2}$, a second-order moment term $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{D}^{2}}$, and a hyperedge $\operatorname{e}^{i2\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D}}$ among modes $\{B,C,D\}$, as shown in Fig. 2(e).
Therefore, in a hypergraph state of $k$-th order, a new hyperedge of $k$-th order can be generated to connect different modes by shearing operations. This effect can also link more pairs in the graph state. In principle, if a hypergraph state contains one $k$-th order hyperedge and another $l$-th order hyperedge which intersect in exactly one mode, shearing operations acting on this shared mode can expand the maximum order to $(k+l-2)$, offering a practical approach to generate more or higher-order hypergraph states with multimode nonlinearity using only local Gaussian operations.
@@ 3.3 Squeezing
Applying a single-mode squeezing operator $S^{a}(s)=\operatorname{e}^{-i\frac{s}{2}(\operatorname{\mathit{\hat{p}}}_{a}\operatorname{\mathit{\hat{q}}}_{a}+\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a})}$ on mode $a$ multiplies the weight of all hyperedges linked to mode $a$ by an additional factor of $\operatorname{e}^{-s}$, while other hyperedges remain unchanged. That is, it updates the edge and weight sets by $(E^{\prime};T^{\prime})=(E;T)\cup(E_{a};\{\tilde{t}\})$, where $\tilde{t}\coloneqq(\operatorname{e}^{-s}-1)\times t_{e}$.
In Fig. 2(f), we show an example by squeezing mode $A$ with $S^{A}(1)$, hence, $t_{AD}=2\operatorname{e}^{-1}$ and $t_{ABC}=\operatorname{e}^{-1}$. Therefore, by using a squeezing operation, the value of weights in the hyperedges can be effectively proportionally adjusted, that is, the strength of the correlation between different modes can be changed.